OpenStudy (anonymous):

Segment AB and BC form a right angle at point B. If A=(-3,-1) and B=(4,4) what is the equation of BC?

OpenStudy (anonymous):

Slope of AB =m \[ m=\frac{4-(-1)}{4-(-3)}=\frac 5 7 \]

OpenStudy (anonymous):

Slope of BC is \[ -\frac 75 \]

OpenStudy (anonymous):

B=(4,4) equation of BC \[ y-4= -\frac 7 5 ( x-4)\\ y=-\frac 7 5 ( x-4)+4 \]

OpenStudy (anonymous):

Im not sure how to do it. These were my choices x+3y=16 -7x-5y=-48 7x-5y=48

OpenStudy (anonymous):

BC would be (4,-1)?????

OpenStudy (anonymous):

-7x-5y=-48

OpenStudy (anonymous):

\[ y-4= -\frac 7 5 ( x-4)\\ 5 y -20 = -7x +28 \\ 7x +5y =48\\ -7x -5y =-48 \] Second choice

OpenStudy (anonymous):

thank you!!!!

OpenStudy (anonymous):

yw